The manager of AllNeeds supermarket wants to estimate µ,, the average amount of money spent per week



for groceries by each student, assuming this amount of money follows approximately a normal distribution.



A random sample of 90 is taken, of which the mean and standard deviation of the amount of money spent



per week for groceries are GHC81.50 and GHC10.36 respectively.



(a) Calculate a 99% confidence interval for µ.



(b) Based on your result in (i), would you say on average a student spends more than GHC30?



(c) How many students must be surveyed if the manager wanted to estimate the true average spending



of students within a margin of .10 with probability .95?

1
Expert's answer
2022-03-19T04:09:06-0400

"\alpha=0.99/2=0.495"

"z_{\alpha/2}=2.575"

"\mu=81.5\plusmn Z_{\alpha/2} \frac{\sigma}{\sqrt{n}}=81.5\plusmn 2.575\frac{10.36}{\sqrt{90}}=81.5 \plusmn 2.8"

b. on average a student spends (78.7-84.3). It is greater then GHC30

c."n=(\frac{Z_{\alpha/2}}{\sigma})^2"

"Z_{\alpha\2}(0.95)=1.96"

"n=(1.96\/0.1)^20.5(1-0.5)=96"